# Cardinality of Real Numbers and Set Axiomatization

### v2

@@ Abstract
The cardinality of real numbers is contingent upon the chosen axiomatic system and the specific models of set theory. According to the diagonal argument within the Zermelo-Fraenkel (ZF) framework, real numbers are non-denumerable. However, their precise cardinality remains a topic of debate, varying with different extensions of the ZF axioms. For example, under the inner-model axiom ‘V-ultimate L’, the Continuum Hypothesis (CH) holds true. In contrast, it is negated if Martin’s Axiom is adopted. Additionally, there is an orthogonal view that assumes there are myriad mathematical universes, some in which the continuum hypothesis is true and others in which it is false, but all equally worth of exploring. This model-dependent viewpoint is underlined by the Löwenheim-Skolem theorem, which asserts that no theory can have only non-denumerable models. Accepting any form of axiomatization implies that terms like ‘denumerable’ and ‘non-denumerable’ are relative. This necessitates the acceptance of the relativization of cardinal numbers, as initially suggested by Skolem and Carnap.
To investigate this uncertainty in cardinality, we analyzed two analogous sets within the frameworks of the ZFC (ZF and Axiom of Choice (AC)) model and Wang’s $\Sigma$ model, which is different from ZFC. The first set, denoted as $S^{\prime}$, comprises all pairs of algebraic numbers, and transcendental numbers inserted between each pair by a specifically defined function $f$. This function places an arbitrarily chosen transcendental number between each algebraic pair. A key attribute of set $S^{\prime}$ in the ZF framework is that it encompasses all possible intervals with algebraic endpoints, defined by these algebraic pairs. This characteristic enables the construction of any Cauchy sequence representing any number using these intervals. By applying the AC, one can systematically identify within set $S^{\prime}$ the intervals constituting any given Cauchy sequence and coherently assemble them to represent any number, even though no individual sequence is explicitly defined within set $S^{\prime}$. Owing to the operational nature of function $f$, every interval identified in these Cauchy sequences is assuredly non-empty. Consequently, in line with the Nested Interval Property (NIP), the intersections of these intervals are also non-empty, containing specific transcendental numbers as defined by the sequences. This leads to the remarkable conclusion that set $S^{\prime}$ is simultaneously complete (any number can be recognized in the set) and countable (a countable number of transcendental numbers is inserted in the countable set).
To expand upon this result, we examined an equivalent set that is also derived by inserting transcendental numbers between algebraic numbers. However, in this case, the transcendental numbers are not selected arbitrarily. Rather, they are systematically constructed and exactly defined by applying diagonal procedures on the sequence of listed numbers. This method is employed to create the layers of Wang’s $\Sigma$ model. In this novel approach, we present the first explicit example that is consistent with that model. The process begins with the sequence of all algebraic numbers representing the 0-th Wang’s layer. The 1-st layer is then constructed through diagonal procedures applied to the 0-th layer. These procedures yield transcendental numbers, which, in conjunction with the 0-th layer, form the first layer. Subsequent layers are similarly created, with each new layer arising from diagonal procedures made on the preceding layer and by adding these transcendental numbers created by these diagonal procedures to the preceding layer, thus progressively constructing Wang’s $\Sigma$ layered structure.
This process is finite and concludes precisely after $\omega$ steps, constrained by the atomic nature of real numbers. Upon examining the properties of the resulting set, and given its equivalence to set $S^{\prime}$, we conclude that real numbers that can be constructively expressed are countable. This finding appears to challenge the traditional diagonal argument. However, it is important to recognize that this model does not encompass all real numbers, particularly those that are non-constructible. Nevertheless, the apparent contradiction persists, as the diagonal argument primarily concerns constructible numbers, since it is applied to well-defined numbers expressed by digital sequences.
The application of the AC was crucial in the construction of set $S^{\prime}$; it allows for the selection of an arbitrary transcendental number for each pair of algebraic numbers. It is also essential in enabling the identification of any interval with algebraic endpoints and the identification of any specific sequence within set $S^{\prime}$, leading to the conclusion that set $S^{\prime}$ is complete and enumerable. While this result might resonate with constructivists who oppose the AC, our further exploration of the diagonal argument revealed elements that conflict with the NIP. This issue is not unique; it also emerges in other proofs of non-denumerability, as detailed in the article. This necessitates a more thorough reevaluation of the countability of real numbers.
@@
Set axioms , Continuum hypothesis , Diagonal argument , Axiom of choice , Denumerability of numbers , V-ultimate L , Cardinal numbers MSC: 11B05
and
@@ 1 Introduction
Gödel conceived a small and constructible model universe called L that is generated by starting with an empty set and iterating it to build bigger and bigger sets. In 1938, he put forward a new axiom for set theory: the axiom V=L. By considering the inner model for set theory in which V=L is true, Gödel was able to prove the relative consistency of the Zermelo-Fraenkel (ZF) axioms and ZF plus the axiom of choice (ZFC), and the generalized Continuum Hypothesis (CH)[1]. Gödel briefly considered proposing to add V=L to the accepted axioms for set theory as a sort of meaning stipulation, but he soon changed his mind. His later view was that V=L is really false, even though it is consistent with set theory if set theory is itself consistent. He believed that CH is not true and that not every uncountable set of reals has the same size as $\mathbb{R}$, and that indeed $2^{\aleph_{0}}=\aleph_{2}$ is true[2, 3]. “One may on good reason suspect that the role of the continuum problem in set theory will be this, that it will finally lead to the discovery of new axioms which will make it possible to disprove Cantor’s conjecture” [5]. Ever since it was established that CH is independent of the ZFC axioms by Gödel’s 1940 proof that the ZFC axioms cannot be used to disprove the continuum hypothesis, and Cohen’s 1963 proof to the contrary that they cannot be used to prove it[4], there have been efforts to add at least one new axiom to ZFC that would illuminate the structure of infinite sets. Currently, the two most promising candidates are the Ultimate-L axiom, or the statement that Ultimate-L is the universe of sets, which is an all-encompassing generalization of Gödel’s model universe of sets with cardinality $\aleph_{1}$; and Martin’s Maximum axiom[6, 7], which asserts that anything conceived of using any forcing construction is a true mathematical entity, so long as the construction satisfies a certain consistency condition, compellingly pointing to $\aleph_{2}$ cardinality.
So, forcing axioms provide an avenue for generic extensions of the universe $V$ of all sets through an iterative process $V\subset V[g_{0}]\subset V[g_{1}]\subset\ldots\subset V[g_{\alpha}]$ of $V$ for any length $\alpha$, in such a way that the final model is itself a generic extension of $V$. Cohen used forcing to enlarge the continuum to include new reals beyond those of the model. But applying the forcing mechanism allows creating arbitrarily many reals, $\aleph_{2}$ or $\aleph_{315}$, etc., a cloud of universes, without the possibility to distinguish which is virtual and which is real, or even whether an object created that way really exists.
To address this ambiguity and establish the actual existence of specific objects rendered possible by Cohen’s method, various forcing axioms are introduced, such as Martin’s Maximum axiom and Woodin’s (*) axiom [8], that allow stating “for all $X$, there exists $Y$, such that $Z$ when referring to properties of sets within the domain.
@@
However, there are inconsistencies between Martin’s and (*) axioms and between these axioms and ZFC [9, 10, 11]. For instance, the statement related to the (*) axiom, “for all sets of $\aleph_{1}$ reals, there exist reals not in those sets,” clearly contradicts the continuum hypothesis. Recent work [12] shows that Martin’s maximum++ (a technical variation of Martin’s maximum) implies Woodin’s (*) axiom, which amalgamates these two prominent axioms of set theory, both implying that there are $\aleph_{2}$ many real numbers.
Yet, there are still open questions. Woodin introduced the axioms (*)+ and (*)++, which apply to the full power set (the set of all subsets) of the reals, strengthening the (*) axiom [11]. But these axioms contradict Martin’s axiom in various models [11], questioning the existential statements about sets of reals in Martin’s maximum framework.
@@
The problem can be further underlined by Löwenheim-Skolem theorem or the intimately related Gödel completeness theorem and its model-theoretic generalizations. It says that a suitable first-order theory (in countable language) has a countable model, that a set can be ‘nondenumerable’ in this relative sense and yet be denumerable ‘in reality’. As Skolem sums it up, “even the notions ‘finite’, ‘infinite’, ‘simply infinite sequence’ and so forth turn out to be merely relative within axiomatic set theory”, which means that the notion of countability is not absolute [13].
@@
One of the consequences of the theorem is the existence of a countable model of set theory, which nevertheless must satisfy the sentence asserting that the real numbers are uncountable. The diagonal argument requires that $\mathbb{R}$ is nondenumerable in all models, that ZF set theory has only nondenumerable models. However, by the Löwenheim-Skolem theorem this is impossible, no theory can have only nondenumerable models. If a theory has a nondenumerable model, it must have denumerably infinite ones as well.
@@
The non-denumerability is in contrast with predicates used in the diagonal argument, available only in a denumerably infinite multitude and yield no more than $\aleph_{0}$ subsets of the set of all integers, i.e., an enumeration of the continuum from outside the axiomatic system [14]. The enumeration from outside utilizes the structure of the system as a whole, which cannot be reached by operations within the system. There are one-to-one correspondences between $\mathbb{R}$ and $\mathbb{N}$, but they all lie outside the given model. It follows that $\mathbb{R}$ is nondenumerable in a relative sense. What is a ‘countable’ set from the point of view of one model may be an uncountable set from the point of view of another model. For any given nondenumerable theory model $M$, exist a countable model $M^{\prime}$ of that same theory restricted to $M$, the totality that the variables of quantification range over ($M$) a proper subset of the $M^{\prime}$. Therefore, it is not possible to rule out denumerable interpretation. Infinite sets may have different cardinal numbers and be nevertheless syntactically of the same cardinal number [14]. In a given axiomatic set theory $M$, there always exists a more comprehensive theory in which all infinite sets of $M$ prove equivalent, namely denumerable. Every non-denumerable set becomes denumerable in a higher system or in an absolute sense.

### v3

# A Countable, Dense, Dedekind-Complete Subset of $\mathbb{R}$ Constructed by Extending $\mathbb{Q}$ via Simultaneous Marking of Closed Intervals with Rational Endpoints

@@ Abstract
This article explores the model-dependent nature of set cardinality, emphasizing that it is not an absolute measure. The cardinality of a set may differ across models, even when those models are internally consistent and complete within their respective universes. This inherent relativity is a foundational concept in modern set theory and underlies why cardinality, and the Continuum Hypothesis (CH), are intrinsically tied to the framework of the chosen model. For example, while the diagonal argument shows that the real numbers are non-denumerable within the Zermelo-Fraenkel (ZF) framework, the precise cardinality of the continuum remains unsettled and varies across different extensions of ZF. Under the inner-model axiom V = Ultimate L, CH holds; in contrast, it is refuted under Martin’s Axiom. There is also a multiverse view, in which the continuum hypothesis is true in some models and false in others, all of which are considered mathematically valid. The Löwenheim–Skolem theorem reinforces this view by showing that no first-order theory can have only non-denumerable models. Thus, notions such as “denumerable” and “non-denumerable” become relative within any axiomatic framework, necessitating the relativization of cardinal numbers, as originally suggested by Skolem and Carnap.
To examine these issues, we construct and analyze two analogous sets: one within the ZFC (ZF plus the Axiom of Choice) framework, and another within Wang’s $\Sigma$ model, which differs from ZFC. Both constructions follow Gödel’s idea of building larger sets from smaller ones.
The first set, $S_{m}$ is constructed from a countable collection $S_{Q}$ of all closed intervals with rational endpoints, arranged in a sequence $\{S_{Q}\}$. All intervals are pre-marked through a simultaneous marking process, which assigned irrational points to all intervals in one unified step, respecting the structure of the sequence $\{S_{Q}\}$ and all its subsequences. Each interval is used exactly once, and no two distinct irrational numbers are assigned to the same interval. While each interval $[a_{n},b_{n}]$ contains infinitely many subintervals $[a^{\prime}_{n},b^{\prime}_{n}]\subset[a_{n},b_{n}]$, each of which is assigned a marked number, the construction ensures that exactly one marked number is assigned to each interval $[a_{n},b_{n}]\in S_{Q}$. All intervals are marked without omission, ensuring that in every interval from the begining, as the outset, is one marked or assigned irrational number. Such marking reults in the set $S_{m}$, which contains all rational endpoints from $S_{Q}$ and the assigned irrational numbers.
Within ZF, $S_{m}$ exhibits two critical properties: it is everywhere dense in $\mathbb{R}$, and as it will be shown it is Dedekind complete. Since $S_{Q}$ contains all closed intervals with rational endpoints, it enables inside $S_{Q}$ identification of any Cauchy sequence representing any real number. The Axiom of Choice plays a critical role here, as it ensures that there exists a global selection function that simultaneously assigns marked numbers to all intervals in $S_{Q}$ respecting the strucute of the sequence $\{S_{Q}\}$ and its subsequences, and for the identification of intervals forming any Cauchy sequence, without explicitly constructing the sequence itself. Because all intervals in such identified sequences are pre-marked from the outset, and that marking respects the nested structure of all intervals, that in all intervals there are numbers simultaneously, ensures that their intersection remains nonempty and includes the irrational number represented by the sequence. This leads to the remarkable conclusion that $S_{m}$ is both complete and countable: a countable set of irrational numbers is distributed across a countable set of intervals, one per interval.
The second set is constructed within Wang’s $\Sigma$ model by systematically inserting irrational numbers between rational numbers using a diagonal procedure. Unlike in $S_{m}$, these irrational numbers are not arbitrarily assigned, but are instead explicitly defined through successive diagonal constructions. The process begins with all rational numbers forming the 00-th layer. The first layer is formed by applying a diagonalization procedure to this layer, thereby generating new irrational numbers. Subsequent layers are built by recursively applying similar diagonal procedures to the union of all previous layers. The process concludes after $\omega$ steps, when further subdivision is no longer possible, consistent with the atomic nature of real numbers. The resulting set is equivalent to $S_{m}$, and supports the conclusion that constructively defined real numbers are countable, apparently challenging the traditional interpretation of the diagonal argument. However, it is important to note that this construction applies only to constructible reals and excludes non-constructible ones. Nevertheless, the contradiction remains significant, as the diagonal argument is typically applied to explicitly definable numbers (i.e., those given by digital sequences).
Further analysis in the paper identifies tensions between foundational principles in set theory, especially between the diagonal argument and the Nested Interval Property (NIP). These issues are not limited to one argument, but recur across multiple classical proofs of uncountability. While these results may resonate with constructivist perspectives, especially those critical of the Axiom of Choice, they also motivate a broader reevaluation of the countability of the real numbers within formal mathematical frameworks.
@@
Set axioms , Continuum hypothesis , Diagonal argument , Denumerability , Dedekind completeness , Wang’s $\Sigma$ model
MSC:
[2020] 11B05
@@ 1 Introduction
Gödel conceived a small and constructible model universe denoted L), generated by iteratively building larger sets from the empty set. In 1938, he put forward a new axiom for set theory: the axiom V=L, asserting that every set is constructible. By considering the inner model in which $\mathsf{V}=\mathsf{L}$ holds, Gödel demonstrated the relative consistency of the Zermelo–Fraenkel axioms (ZF), ZF augmented with the Axiom of Choice (ZFC), and the Generalized Continuum Hypothesis (CH) [1]. While he initially entertained the idea of adopting $\mathsf{V}=\mathsf{L}$ as a foundational axiom, he later rejected it, believing the axiom to be false despite its consistency within ZF. In his later view, the continuum hypothesis itself was not true, and not every uncountable set of real numbers has the same cardinality as $\mathbb{R}$. He conjectured instead that $2^{\aleph_{0}}=\aleph_{2}$ [2, 3]. As he wrote, “One may on good reason suspect that the role of the continuum problem in set theory will be this: that it will finally lead to the discovery of new axioms which will make it possible to disprove Cantor’s conjecture” [4].
The independence of CH from ZFC was established through Gödel’s 1940 proof that CH cannot be disproven from the ZFC axioms, followed by Cohen’s 1963 result that it also cannot be proven within them [5]. This led to ongoing efforts to introduce new axioms to ZFC that might resolve the continuum problem. Two of the most prominent candidates are the Ultimate-$\mathsf{L}$ axiom, which is an all-encompassing generalization of Gödel’s constructible universe consistent with CH and the cardinality $\aleph_{1}$; and Martin’s Maximum axiom[6, 7], which asserts that anything conceived of using any forcing construction is a true mathematical entity, so long as the construction satisfies a certain consistency condition, compellingly pointing to $\aleph_{2}$ cardinality.
Forcing axioms provide a method for extending the universe $V$ of all sets through an iterative process of generic extensions:
$V\subset V[g_{0}]\subset V[g_{1}]\subset\cdots\subset V[g_{\alpha}],$ (1)
where each $V[g_{\alpha}]$ represents a new universe generated by forcing over the previous one. Cohen’s method of forcing allows the creation of arbitrarily many new reals, beyond those of the model, resulting in extensions where the continuum has cardinality $\aleph_{2},\aleph_{315}$, or other transfinite values. However, this procedure produces a multiplicity of universes, a cloud of univrses, each with a different structure of the continuum, without a canonical way to distinguish which is virtual and which is real, or even if any reflects the actual existence of mathematical objects.
To address this ambiguity, additional forcing axioms such as Martin’s Maximum and Woodin’s ()(*)( ∗ ) axiom have been introduced [8]. These axioms enable the formulation of statements like “for all $X$, there exists $Y$ such that $Z$ when referring to properties within a universe of sets, thus promoting certain existential principles that guide set-theoretic constructions.
@@
However, there are inconsistencies between Martin’s and (*) axioms and between these axioms and ZFC [9, 10, 11]. For example, the statement entailed by ()(*)( ∗ ), “for all sets of $\aleph_{1}$ reals, there exist reals not in those sets”, clearly contradicts CH. Recent work by Aspero and Schindler [12] shows that a strong version, Martin’s Maximum++, implies Woodin’s ()(*)( ∗ ), further linking two of the most influential axioms. Both imply that the continuum has cardinality $\aleph_{2}$, yet both also highlight unresolved tensions within the set-theoretic landscape.
Woodin has proposed further axioms $(*)^{+}$ and $(*)^{++}$, which quantify over the entire power set of the reals, extending the expressive strength of ()(*)( ∗ ). However, these stronger axioms are known to conflict with Martin’s Maximum in certain models [11], questioning the existential statements about sets of reals in Martin’s maximum axiom, and if if the reals can be coherently maintained in a unified set-theoretic framework.
@@
The problem can be further underlined by Löwenheim-Skolem theorem or the intimately related Gödel completeness theorem and its model-theoretic generalizations. These results imply that any first-order theory formulated in a countable language must have a countable model, that a set can be ‘nondenumerable’ in this relative sense and yet be denumerable ‘in reality’. As Skolem sums it up, “even the notions ‘finite’, ‘infinite’, ‘simply infinite sequence’ and so forth turn out to be merely relative within axiomatic set theory” [13]. Thus, the notion of countability is not absolute, but model-relative.
@@
One of the consequences of the theorem is the existence of a countable model of set theory, which nevertheless must satisfy the sentence asserting that the real numbers are uncountable. The diagonal argument requires that $\mathbb{R}$ is nondenumerable in all models, that ZF set theory has only nondenumerable models. However, by the Löwenheim-Skolem theorem this is impossible, no theory can have only nondenumerable models. If a theory has a nondenumerable model, it must have denumerably infinite ones as well. Any model of ZFC that contains uncountable sets must admit a countable model (from the perspective of a larger meta-theory) that satisfies the same first-order properties.
@@
The non-denumerability is in contrast with predicates used in the diagonal argument, available only in a denumerably infinite multitude and yield no more than $\aleph_{0}$ subsets of the set of all integers, i.e., an enumeration of the continuum from outside the axiomatic system [14]. The enumeration from outside utilizes the structure of the system as a whole, which cannot be reached by operations within the system. There are one-to-one correspondences between $\mathbb{R}$ and $\mathbb{N}$, but they all lie outside the given model.
Consequently, the uncountability of the reals becomes a model-relative property. A set that is countable in one model may be uncountable in another. For every non-denumerable model $M$ of set theory, there exists a countable model $M^{\prime}\subset M$ satisfying the same first-order axioms. Variables in $M^{\prime}$ range over a proper subset of those in $M$, allowing a denumerable interpretation of otherwise uncountable sets. Therefore, it is not possible to rule out denumerable interpretation. Infinite sets may have different cardinal numbers and be nevertheless syntactically of the same cardinal number [14]. In a given axiomatic set theory $M$, there always exists a more comprehensive theory in which all infinite sets of $M$ prove equivalent, namely denumerable. Every non-denumerable set becomes denumerable in a higher system or in an absolute sense.